# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ IV.1 Graph families for Laplacian matrix-based systems
We survey 30 graph families and their Laplacian matrices as a part of our LSP, with the details provided in Table 2. Of them, we find 1 best and 1 better graph family, along with 5 good graph families. The rest of the families are found to be bad. We present our results in detail for the former two classes of graphs as they are of most interest for quantum advantage. Results the latter two classes can be found in Figs. A.8 and A.9 of the Appendix.
IV.1.1 Best graph families
As Table 2 indicates, the hypercube graph family is the only best graph family. Fig. 2 presents the relevant data. Both $\kappa$ and $s$ grow as log(N)(N), and $N\sim 2^{n}$. Among the non-tree graph families, this is the only graph family from our candidates that is characterized by an exponential system size growth. The exponential growth of system size is a critical requirement for a graph to be a best graph, over and above $\kappa$ and $s$ scaling at most polylogarithmically. We also note from Fig. 2(b) that the CKS(1) and the AQC(1) algorithms offer quantum advantage, that is, their curves cross the R(N)=1R({N})=1 line, well before the other approaches do, whereas HHL does not show advantage yet over the sampled range of $N$ values. Fig. 2(c) presents the advantage offered by HHL algorithm with respect to CLS in the extrapolated data regime. In evaluating the runtime ratio $\tilde{R}(n)$, we exclude all the pre-factors. We emphasize that our analysis is purely based on graph theoretic considerations and neglects implementation overheads such as the number of physical qubits given a quantum error correcting code, cost of magic state distillation, etc. Thus, our numerical results for the crossover point cannot be taken quite literally but rather as qualitative runtime performance indicators even at small system sizes. We adopt this perspective in discussing the crossover points for the subsequent graph families.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
Figure 5: Better graph families of incidence-matrix based systems: The first, second and third column of each panel represents $\kappa$, $s$ vs. system size $N^{\prime}=N+M$, $R(N^{\prime})$ vs $N^{\prime}$ and $\tilde{R}(n)$ vs. $n$ accompanied by a reference linear curve, $n$, for the first three better graph families listed in Table 3.
IV.1.2 Better graph families
Fig. 3 shows the better graph identified in our survey: the modified Margulis-Gabber-Galil graph family. The Margulis-Gabber-Galil construction is modified to avoid self loops and parallel edges. As Table 2 shows, while $\kappa$ and $s$ both grow at most polylogarithmically, which is desirable, the system size itself grows only polynomially, resulting in a net polynomial advantage with these two graphs. For the modified Margulis-Gabber-Galil graph family, Fig. 3(b) shows that Psi-HHL, CKS(1), and AQC(1) algorithms already begin above the $R(N)=1$ line while HHL crosses over towards the end of our sampled range.
IV.1.3 Good graph families
Fig. A.8 of the Appendix presents our plots for the good graphs that we find. As Table 2 shows, we only get a sub-linear advantage with these graphs even though $\kappa$ and $s$ scale at most polylogarithmically because their system size grows linearly (with the exception of the Sudoku graph, where the system size grows quartic in $n$ but $s$ grows as $n^{2}$). In other words, $\tilde{R}$ has only an $n$ to compete with the accompanying $ll(n)/\mathrm{polylog}(n)$ term, and therefore the net advantage is only sub-linear. For the same reason, none of the random graphs considered in our study gives an exponential or a polynomial advantage, but only at most a sub-linear advantage. We emphasize that this conclusion depends on our choice of input parameters for the random graphs, and therefore alternative parameter regimes may yield better runtime performance. We also add that for the cases where system size growth is linear, one could restrict $n$ values for which a graph is allowed via additional conditions in the construction or find an application where such networks occur only for certain $N$ values, but we do not consider those cases in our study.
IV.1.4 Bad graph families
Fig. A.9 of the Appendix gives our results for bad graphs. For all of them, either $\kappa$ or $s$ grow at least polynomially. This, in conjunction with the system size growing linearly (with the exception of tree graphs, which we shall comment on soon), leads to no advantage with HHL.
It is interesting to note that with CKS(1) and AQC(1) algorithms, some of these bad graphs graduate to the good graph category. This still requires the product of $\kappa$ or $s$ to be less than quadratic for those graphs. Our results also show that the dream QLS further improves over CKS(1)/AQC(1) on bad graph families, and offers a polynomial advantage, thus underscoring the need for further work on the algorithmic sector on QLSs.
@@ IV.1.4 Bad graph families
Of special interest among the families in this category is the transition that we observe in the growth of condition number for grid 2d, hexagonal lattice and triangular lattice graph families. In all the three cases, $\kappa$ grows slowly up to a certain system size and beyond it, grows polynomially. Since we are concerned with $\kappa$ scaling at large system sizes to calculate runtime ratios, we consider the functional form of $\kappa$ to be polynomial in $N$. Although sparsity remains constant for these graph families, they provide no advantage with any QLS because of their polynomial growth in $\kappa$.
We now move to the three tree graph families in this category. The system size and $\kappa$ grow exponentially, and hence these graphs yield no advantage even with sparsity being a constant or a polylog function. We observe the same behaviour with a variant of grid 2d graph, where the number of vertices along a row and column are the same and grow exponentially, denoted by ‘Grid 2d (r=c)’ in Table 2. However, interestingly, with CKS(1)/AQC(1) as well as the dream QLS approaches, we find that these graph families offer an exponential advantage, with the crucial catch that the runtime complexity with these quantum algorithms is still exponential and thus the relative exponential speed-up has no practical relevance. We term this observed behaviour as examples of futile exponential advantage. We also add at this juncture that we can generalize and state that balanced $r$-ary tree graph families, where $r\in\mathbb{Z}^{+}$, have infinite such families that all offer futile exponential advantage.
Table 3: Table presenting data on incidence matrices of directed graph families that have been considered for this study.
@@
Graph Graph name $N^{\prime}$ $\kappa$ $s$ $t_{\mathrm{HHL}}(n)$ $t_{\mathrm{CLS}}(n)$ Advantage with HHL, Advantage
type (Source and sink vertices (Yes/No)) $\tilde{R}(n)=t_{\mathrm{CLS}}(n)/t_{\mathrm{HHL}}(n)$ with CKS(1)/AQC(1),
with dream
QLS
Best Directed Hypercube $^{n} n2n^{2}$ $n$ $n^{10}$ $2^{n}n^{3}\mathrm{log}(n)$ exp, exp, exp
graph (Yes) $2^{n}\mathrm{log}(n)/n^{7}$
Better Gaussian random partition (No) $n^{2}$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n^{2}\mathrm{log}^{9/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{25/2}(n)$ poly, poly
Gaussian random partition (Yes) $n^{2}$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n^{2}\mathrm{log}^{9/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{25/2}(n)$ poly, poly
Planted partition graph (No) $n^{2}$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n^{2}\mathrm{log}^{9/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{25/2}(n)$ poly, poly
Planted partition graph (Yes) $n^{2}$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n^{2}\mathrm{log}^{9/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{25/2}(n)$ poly, poly
Navigable small world graph (No) $n^{2}$ $\mathrm{log}^{3}(n)$ $\mathrm{log}(n)$ $\mathrm{log}^{13}(n)$ $n^{2}\mathrm{log}^{5/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{21/2}(n)$ poly, poly
Navigable small world graph (Yes) $n^{2}$ $\mathrm{log}^{3}(n)$ $\mathrm{log}(n)$ $\mathrm{log}^{13}(n)$ $n^{2}\mathrm{log}^{5/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{21/2}(n)$ poly, poly
Gn,p graph (No) $n^{2}$ $\mathrm{log}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{7/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{15/2}(n)$ poly, poly
Gn,p graph (Yes) $n^{2}$ $\mathrm{log}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{7/2}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{15/2}(n)$ poly, poly
Paley graph (No) $n^{3}$ $c$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{8}(n)$ $n^{3}\mathrm{log}^{3}(n)ll(n)$ poly, $n^{3}ll(n)/\mathrm{log}^{5}(n)$ poly, poly
Good Random uniform $n$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n\mathrm{log}^{9/2}(n)ll(n)$ sub-linear, sub-linear,
k-out graph (No) $nll(n)/\mathrm{log}^{25/2}(n)$ sub-linear
Random uniform $n$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n\mathrm{log}^{9/2}(n)ll(n)$ sub-linear, sub-linear,
k-out graph (Yes) $nll(n)/\mathrm{log}^{25/2}(n)$ sub-linear
Scale-free graph (No) $n$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n\mathrm{log}^{9/2}(n)ll(n)$ sub-linear, sub-linear,
$nll(n)/\mathrm{log}^{25/2}(n)$ sub-linear
Scale-free graph (Yes) $n$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n\mathrm{log}^{9/2}(n)ll(n)$ sub-linear, sub-linear,
$nll(n)/\mathrm{log}^{25/2}(n)$ sub-linear
Bad GN graph(No) $n$ $^{2} nn$ $n^{8}\mathrm{log}^{2}(n)$ $n^{3}ll(n)$ none, $ll(n)/n^{5}\mathrm{log}^{2}(n)$ none,
sub-linear
GN graph(Yes) $n$ $n^{2}$ $\mathrm{log}^{3}(n)$ $n^{6}\mathrm{log}^{8}(n)$ $n^{2}\mathrm{log}^{3}(n)ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{5}(n)$ none,
sub-linear
GNC graph(No) $n^{2}$ $\mathrm{log}^{3}(n)$ $n^{4}$ $n^{8}\mathrm{log}^{11}(n)$ $n^{6}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/n^{2}\mathrm{log}^{19/2}(n)$ poly, poly
GNC graph(Yes) $n^{2}$ $\mathrm{log}^{3}(n)$ $n^{4}$ $n^{8}\mathrm{log}^{11}(n)$ $n^{6}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/n^{2}\mathrm{log}^{19/2}(n)$ poly, poly
GNR graph(No) $n$ $\mathrm{log}^{3}(n)$ $n$ $n^{2}\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{19/2}(n)$ sub-linear,
poly
GNR graph(Yes) $n$ $n^{2}$ $\mathrm{log}^{3}(n)$ $n^{6}\mathrm{log}^{8}(n)$ $n^{2}\mathrm{log}^{3}(n)ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{5}(n)$ none,
sub-linear

### v4

@@ IV.1 Graph families for Laplacian matrix-based systems
We survey 30 graph families and their Laplacian matrices as a part of our study, with the details provided in Tables 2 and 3. We find 7 good and 23 bad graph families. We present two instances each for good and bad graph categories in Figs. 2 and 3. We refer the reader to Figs. S.8 and S.9 of the Supplemental Material for the rest of the good and bad graph families respectively.
IV.1.1 Good graph families
As Table 2 indicates, all the graph families in this category show at most polylogarithmic growth in $\kappa$ and $s$ (with the exception of Sudoku graph, whose sparsity function grows as $\sqrt{N}$), leading to an exponential advantage with HHL. The exponential advantage is retained with CKS(1), AQC(1) and dream QLS algorithms. Furthermore, with CKS(1) and AQC(1) algorithms, good graph families offer advantage well before the other algorithms do in $N$. This is witnessed in the plot of $R(N)$ vs $N$, where the regime of advantage starts when the curve crosses R(N)=1R({N})=1 line. Among the good graph families that we identify, the admissible $N$ values in the hypercube graph family grow as $N=2^{n}$, and thus the problem is classically hard. As mentioned earlier, our analyses does not account for practical overheads in estimating quantum advantage. Thus, our numerical results for the crossover point cannot be taken quite literally but rather as qualitative runtime performance indicators even at small system sizes. We adopt this perspective in discussing the crossover points for the subsequent graph families.
IV.1.2 Bad graph families
From Table 3, we see that either $\kappa$ or $s$ grow at least polynomially resulting in no advantage with the HHL algorithm. It is interesting to note that with CKS(1) and AQC(1) algorithms, 15 graph families graduate to good graph family and yield exponential advantage, while with dream QLS, 16 graph families offer exponential advantage.
@@ IV.1.2 Bad graph families
Of special interest among the families in this category is the transition that we observe in the growth of condition number for grid 2d, hexagonal lattice and triangular lattice graph families. For all the three cases, $\kappa$ grows slowly up to a certain system size and beyond it, grows polynomially. Since we are concerned with $\kappa$ scaling at large system sizes to calculate runtime ratios, we consider the functional form of $\kappa$ to be polynomial in $N$. Although sparsity remains constant for these graph families, they provide no advantage with HHL because of their polynomial growth in $\kappa$.
[…]
Graph name $\kappa$ $s$ $t_{\mathrm{HHL}}$ $t_{\mathrm{CKS(1)}}(\text{or }t_{\mathrm{AQC(1)}})$ $t_{\mathrm{dream\ solver}}$ $t_{\mathrm{CLS}}$
(Source and sink vertices (Yes/No))
Directed Hypercube $\mathrm{log}^{2}(N^{\prime})$ $\mathrm{log}(N^{\prime})$ $\mathrm{log}^{10}(N^{\prime})$ $\mathrm{log}^{4}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{7/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{2}(N^{\prime})ll(N^{\prime})$
graph (Yes)
Gaussian random $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
partition (No)
Gaussian random $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
partition (Yes)
Planted partition $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
graph (No)
Planted partition $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
graph (Yes)
Navigable small $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}(N^{\prime})$ $\mathrm{log}^{13}(N^{\prime})$ $\mathrm{log}^{5}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{5/2}(N^{\prime})ll(N^{\prime})$
world graph (No)
Navigable small $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}(N^{\prime})$ $\mathrm{log}^{13}(N^{\prime})$ $\mathrm{log}^{5}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{5/2}(N^{\prime})ll(N^{\prime})$
world graph (Yes)
Gn,p graph (No) $\mathrm{log}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{11}(N^{\prime})$ $\mathrm{log}^{5}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{7/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{7/2}(N^{\prime})ll(N^{\prime})$
Gn,p graph (Yes) $\mathrm{log}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{11}(N^{\prime})$ $\mathrm{log}^{5}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{7/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{7/2}(N^{\prime})ll(N^{\prime})$
Paley graph (No) $c$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{8}(N^{\prime})$ $\mathrm{log}^{4}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{5/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{3}(N^{\prime})ll(N^{\prime})$
Random uniform $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
k-out graph (No)
Random uniform $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
k-out graph (Yes)
Scale-free graph (No) $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
Scale-free graph (Yes) $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{3}(N^{\prime})$ $\mathrm{log}^{17}(N^{\prime})$ $\mathrm{log}^{7}(N^{\prime})ll(N^{\prime})$ $\mathrm{log}^{11/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime}\mathrm{log}^{9/2}(N^{\prime})ll(N^{\prime})$
Table 5: Table presenting our data for incidence matrix-based graph families that offer no advantage with the HHL algorithm. We also give the advantage offered by these graph families with CKS(1) (or AQC(1)) and dream QLS algorithms (see Eqns. 2 and 3 respectively). Here, ‘Nil’ represents no advantage and ‘Exp’ represents exponential advantage in the table.
@@
Graph name $\kappa$ $s$ $t_{\mathrm{HHL}}$ $t_{\mathrm{CKS(1)}}(\text{or }t_{\mathrm{AQC(1)}}),$ $t_{\mathrm{dream\ solver}}$ $t_{\mathrm{CLS}}$ Advantage with
(Random (Yes/No)) HHL,
CKS(1)(or AQC(1)),
dream solver
GN graph(No) $^{\prime 2} NN^{\prime}$ $N^{\prime 8}\mathrm{log}^{2}(N^{\prime})$ $N^{\prime 3}\mathrm{log}(N^{\prime})$ $N^{\prime 5/2}\mathrm{log}(N^{\prime})ll(N^{\prime})$ $N^{\prime 3}ll(N^{\prime})$ Nil, Nil, Exp
$\mathrm{log}(N^{\prime 3}\mathrm{log}(N^{\prime}))$
GN graph(Yes) $N^{\prime 2}$ $\mathrm{log}^{3}(N^{\prime})$ $N^{\prime 6}\mathrm{log}^{8}(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{4}(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{5/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{3}(N^{\prime})ll(N^{\prime})$ Nil, Nil, Nil
$\mathrm{log}(N^{\prime 2}\mathrm{log}^{4}(N^{\prime}))$
GNC graph(No) $\mathrm{log}^{3}(N^{\prime})$ $N^{\prime 2}$ $N^{\prime 4}\mathrm{log}^{11}(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{4}(N^{\prime})$ $N^{\prime}\mathrm{log}^{4}(N^{\prime})ll(N^{\prime})$ $N^{\prime 3}\mathrm{log}^{3/2}(N^{\prime})ll(N^{\prime})$ Nil, Exp, Exp
$\mathrm{log}(N^{\prime 2}\mathrm{log}^{4}(N^{\prime}))$
GNC graph(Yes) $\mathrm{log}^{3}(N^{\prime})$ $N^{\prime 2}$ $N^{\prime 4}\mathrm{log}^{11}(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{4}(N^{\prime})$ $N^{\prime}\mathrm{log}^{4}(N^{\prime})ll(N^{\prime})$ $N^{\prime 3}\mathrm{log}^{3/2}(N^{\prime})ll(N^{\prime})$ Nil, Exp, Exp
$\mathrm{log}(N^{\prime 2}\mathrm{log}^{4}(N^{\prime}))$
GNR graph(No) $\mathrm{log}^{3}(N^{\prime})$ $N^{\prime}$ $N^{\prime 2}\mathrm{log}^{11}(N^{\prime})$ $N^{\prime}\mathrm{log}^{4}(N^{\prime})$ $\sqrt{N^{\prime}}\mathrm{log}^{4}(N^{\prime})ll(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{3/2}(N^{\prime})ll(N^{\prime})$ Nil, Exp, Exp
$\mathrm{log}(N^{\prime}\mathrm{log}^{4}(N^{\prime}))$
GNR graph(Yes) $N^{\prime 2}$ $\mathrm{log}^{3}(N^{\prime})$ $N^{\prime 6}\mathrm{log}^{8}(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{4}(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{5/2}(N^{\prime})ll(N^{\prime})$ $N^{\prime 2}\mathrm{log}^{3}(N^{\prime})ll(N^{\prime})$ Nil, Nil, Nil
$\mathrm{log}(N^{\prime 2}\mathrm{log}^{4}(N^{\prime}))$